### v1

# Model-Based Soft Maximization of Suitable Metrics of Long-Term Human Power

Abstract
Power is a key concept in AI safety: power-seeking as an instrumental goal, sudden or gradual disempowerment of humans, power balance in human-AI interaction and international AI governance. At the same time, power as the ability to pursue diverse goals is essential for wellbeing.
This paper explores the idea of promoting both safety and wellbeing by forcing AI agents explicitly to empower humans and to manage the power balance between humans and AI agents in a desirable way. Using a principled, partially axiomatic approach, we design a parametrizable and decomposable objective function that represents an inequality- and risk-averse long-term aggregate of human power. It takes into account humans’ bounded rationality and social norms, and, crucially, considers a wide variety of possible human goals.
We derive algorithms for computing that metric by backward induction or approximating it via a form of multi-agent reinforcement learning from a given world model. We exemplify the consequences of (softly) maximizing this metric in a variety of paradigmatic situations and describe what instrumental sub-goals it will likely imply. Our cautious assessment is that softly maximizing suitable aggregate metrics of human power might constitute a beneficial objective for agentic AI systems that is safer than direct utility-based objectives.
1 Introduction
A duty to empower others, especially those with little power, can be defended on consequentialist (Sen 2014), deontological (Hill Jr 2002), and virtue ethical Nussbaum (2019) grounds. At the same time, gradual or sudden disempowerment of humans due to misaligned A(G)I, e.g. due to power-seeking (Turner et al. 2019) or non-corrigible over-optimization of misaligned reward functions (Gao, Schulman, and Hilton 2023), is a key AI safety risk.
Accordingly, some papers explored tasking AI agents explicitly with the empowerment of individual humans, mainly inspired by the information-theoretic channel capacity $\cal E$ between human actions and environmental states (called ‘empowerment’ in (Klyubin, Polani, and Nehaniv 2005)). As that metric is hard to compute, Du et al. (2020) use a proxy based on the variance of terminal states reached under a random policy, while Myers et al. (2024) use contrastive learning of latent state representations to approximate $\cal E$. Salge and Polani (2017) discuss the ‘empowerment’-based approach in more depth for the single-human case.
This paper extends the theoretical foundations of this approach in two ways. We develop an alternative metric of human power that is directly based on the key aspect of power—the ability to attain a wide range of possible goals—and that explicitly and transparently incorporates humans’ knowledge about the state of the world and the actions of the AI agent, their expectations about others’ behavior, e.g. due to social norms, and their bounded rationality. And, using a desiderata-guided approach similar to social choice and welfare theory, we develop an objective function for an AI agent interacting with populations of humans, based on an aggregate of that metric across humans and time that gives the agent desirable incentives such as communicating well (Reddy, Levine, and Dragan 2022), following orders, being corrigible (Potham and Harms 2025), avoiding irreversible changes in the environment, protecting humans and itself from harm and disempowerment, allocating resources fairly and sustainably, and acting “appropriately” by following relevant social norms (Leibo et al. 2024).
Though based on possible human goals, our approach avoids trying to learn individuals’ actual, current goals, because human preferences are changing and non-identifiable (Cao, Cohen, and Szpruch 2021; Banerjee and Duflo 2011) and their prediction unavoidably uncertain (Baker, Saxe, and Tenenbaum 2011). Some critics of a preference-based approach argue for a values-based approach instead (Lowe et al. 2025), which however requires an even more semantic world understanding by the AI agent. A deep semantic understanding is also required in the ‘freedoms’-based conception of AI ethics in (London and Heidari 2024), and compiling its required list of ‘fundamental capabilities’ is difficult (Robeyns 2006). By contrast, like the ‘empowerment’-based approach, our metrics are mostly based on a structural understanding of possibly dynamics, interactions, and transition probabilities and aim to avoid semantic issues by relating human power to the ability to bring about just any possible conditions they might happen to desire.
For theoretical convenience, we work in a model-based setting where the AI agent can plan on the basis of a decent stochastic world model like the “scientist AI” envisioned in Bengio et al. (2025). After developing our human power metric in Section 2, we shortly describe algorithms for softly maximizing it in Section 3, before reporting first insights about the resulting behavior in Section 4.
2 Measuring, Aggregating, and Maximizing Human (Em)Power(ment)
Figure 1: Overview of proposed approach for deriving power-managing policies for a general-purpose AGI system
Framework
We assume a robot $r$ interacting with several humans $h\in\mathcal{H}$. The robot models the world as a (fully or partially observed) stochastic game with states $s\in\mathcal{S}$ between itself, who gets no extrinsic reward, and the humans who have many possible and potentially changing goals $g_{h}\in\mathcal{G}_{h}$ and get goal-dependent rewards $U_{h}(s^{\prime},g_{h})$ , discounted at factors $\gamma_{h}<1$. Action sets $\mathcal{A}_{r}(s),\mathcal{A}_{h}(s)$ may be state-dependent, action profiles are written $a=(a_{r},a_{\mathcal{H}})=(a_{r},a_{h},a_{-h})$, and the transition kernel is $P(s^{\prime}|s,a)$.
Crucially, we assume the robot neither knows nor forms beliefs about the actual goals of the humans. It does make assumptions about their level of rationality and their beliefs about each others’ behavior, reflected in additional model parameters $\nu_{h}$, $\pi^{0}_{h}$, $\beta_{h}$, and $\mu_{-h}$, as detailed below.
Our task is to design an algorithm for computing a policy $\pi_{r}$ for the robot that implements the aggregate human power maximization objective. For tractability, we formalize this objective as an expected discounted return, $V_{r}=\operatorname*{\mathbb{E}}_{s_{\geq 0}}\sum_{t\geq 0}\gamma_{r}^{t}U_{r}(s_{t})$, based on an intrinsic reward $U_{r}(s)$ representing a suitable assessment of the aggregate human power in $s$. To design $U_{r}(s)$, we first design a power metric $W_{h}(s)$ for individual humans, and then an aggregation function $U_{r}(s)=F_{U}((W_{h}(s))_{h\in\mathcal{H}})$. In designing $W_{h}$ and $F_{U}$, we have several objectives relating to their tractability, interpretability, and the behavioral incentives that they give the robot, in particular relating to AI safety and ethics.
Desideratum Corresponding design choice or world model requirement
$V_{h}$ recursively computable Define via Bellman equation: $V_{h}(s,g_{h})=\operatorname*{\mathbb{E}}_{s^{\prime}}(U_{h}(s^{\prime},g_{h})+\gamma_{h}V_{h}(s^{\prime},g_{h}))$
$V_{h}$ comparable across humans Assume goals $g_{h}$ are events: $g_{h}\subseteq\mathcal{S}$, $U_{h}(s^{\prime},g_{h})=1_{s^{\prime}\in g_{h}}$
$V_{h}$ has natural interpretation Make $V_{h}$ a goal-reaching probability: $s\in g_{h}$, $s^{\prime}$ reachable from $s\Rightarrow s^{\prime}\notin g_{h}$
$r$ incentivized to make commitments Base $\pi_{h}$ on $h$ assuming $r$ takes worst action not ruled out by $r$’s commitments
$r$ considers $h$’s bounded rationality Base $\pi_{h}$ on varying, unstable goals, habits, social norms, mutual expectations
$W_{h}$ based on $r$’s best estimate of $V_{h}$ Distinguish $h$’s simulated estimate $V^{m}_{h}$ for $\pi_{h}$ and $r$’s estimate $V^{e}_{h}$ for $W_{h}$
$W_{h}$ indep. of “unaffected” goals Use separable ansatz: $W_{h}=F^{G}(\sum_{g_{h}}f^{G}(V^{e}_{h}))$ ($\Rightarrow$ enables stoch. approx.)
$W_{h}$ additive across indep. subgames $F^{G}=\log_{2}$ ($\Rightarrow W_{h}=$ “certainty-equivalent” effective no. of binary choices)
$W_{h}>-\infty$ Make set $\mathcal{G}_{h}$ of possible goals wide enough to cover all possible trajectories
Range of $W_{h}$ is symmetric around 0 Make each trajectory fulfill exactly one $g_{h}\in\mathcal{G}_{h}$ and put $\zeta=2$ (see below)
$U_{r}$ indep. of “unconcerned” agents Use separable ansatz: $U_{r}=F^{H}(\sum_{h}f^{H}(W_{h}))$ ($\Rightarrow$ enables stoch. approx.)
Pigou–Dalton-type inequality aversion Make $f^{H}$ strictly concave, e.g., $f^{H}(w)=-2^{-\xi w}$ with $\xi>0$
Protect a human’s “last” bit of power Choose $\xi\geq 1$, e.g., $\xi=1$
$r$ cares for current and later human power Base $\pi_{r}$ on $V_{r}(s_{0})=\operatorname*{\mathbb{E}}_{s_{\geq 0}}\sum_{t=0}^{\infty}\gamma_{r}^{t}U_{r}(s_{t})$, not just $U_{r}(s_{0})$
Limit intertemporal power trading Also make $F^{H}$ strictly concave (intertemporal inequality aversion)
$\pi_{r}$ indep. of common rescaling of $V^{e}_{h}$ Use power laws: $f^{G}(v)=v^{\zeta}$, $\zeta>0$; $f^{H}(w)=-2^{-\xi w}$, $\xi\geq 1$;
$F^{H}(y)=-(-y)^{\eta}$, $\eta>1$; and $\pi(s)(a)\propto(-Q_{r}(s,a_{r}))^{-\beta_{r}}$, $0\leq\beta_{r}\leq\infty$
$r$ incentivized to reduce uncertainty Choose $\zeta>1$ (risk aversion, preference for reliability)
Avoid risks from over-optimization Choose $\beta_{r}<\infty$ (soft optimization, exploration)
Table 1: Desiderata and corresponding metric design choices for metrics of humans’ goal-attainment ability $V_{h}(s,g_{h})$, momentary individual power $W_{h}(s)$, momentary aggregate power $U_{r}(s)$, long-term total human power $V_{r}(s)$ for soft maximization by an AGI system (“robot”) $r$, $r$’s prior on human behavior $\pi_{h}$ used to estimate $V_{r}$, and its own resulting policy $\pi_{r}$.
2.1 Individual-Level Metric: ICCEA Power
Our power metric basically measures how many goals a human can achieve. It’s built in three steps: first, defining what it means to achieve a goal; second, adjusting for the human’s own cognitive limits; and third, aggregating this ability across all possible goals into a single number. Rather than trying to capture the full range of subtle aspects of existing notions of human power, we focus on those aspects we believe a robot can robustly infer from the structure of its world model, encoded in state and action sets, transition kernel, and observation functions. Since we want to incentivize the robot to remove constraints and uncertainties, share information, make commitments, and improve human cognition, coordination, and cooperation, our power metric will depend to some degree on $r$’s model of human decision making.
We start with the intuition that humans have ‘goals’ and that ‘power’ is about the means to achieve a wide range of goals. We aim to measure informationally and cognitively constrained effective autonomous (ICCEA) power: essentially how many goals a human can freely choose to reach with more or less certainty, given their information, cognitive capabilities, and others’ behavior. To design the ICCEA power metric, we start with the ability to reach specific goals, then aggregate this over goals, humans, and finally time and uncertainty in ways that have desirable properties.
Goals
As an easily interpretable compromise between the reachability (Krakovna et al. 2018) and attainable utility (Turner, Hadfield-Menell, and Tadepalli 2020) approaches, but deviating significantly from the information-theoretic approach (Klyubin, Polani, and Nehaniv 2005) that does not explicitly involve any goals, we assume that each possible temporary goal $g_{h}\in G_{h}$ that $h\in\mathcal{H}$ might have is to reach some set of states, $g_{h}\subseteq\mathcal{S}$, representing a possibly desirable event, and that the corresponding reward function is the indicator function $U_{h}(s,g_{h})=1_{s\in g_{h}}$. In other words, $\mathcal{G}_{h}\subseteq{\cal P}(\mathcal{S})$, the power set of $\mathcal{S}$. We require that the elements of $g_{h}$ are mutually unreachable (e.g., are all terminal states or states for a particular time-point $t$), so that the resulting value $V_{h}^{\pi}(s)$ is simply the probability of the desirable event happening under policy profile $\pi$. This grounds $h$’s goal attainment or ‘attainable utility’ in probability, bounded within [0,1][0,1][ 0 , 1 ], which avoids problems with interpersonal utility comparison such as dominance by “utility monsters”. To avoid problems around probability zero, we also require that $G_{h}$ is wide enough so that each possible state trajectory makes some of $h$’s possible goals fulfilled. [^1] Otherwise $X_{h}^{-\xi}$ in eqn. (10) could get infinite. We believe that restricting the model to this very basic type of goal will simplify the derivation of $G_{h}$ from learned latent representations of generic world states and generic human goals as encoded in language or foundation models, will obviate the need for individual-level data about a particular human’s possible goals, and will thus mitigate risks arising from misaligned models of possible human goals. [^2] An even more restrictive choice would be to identify goals with single states, but this seem implausibly specific in complex, partially observed, multi-agent environments, and it would lead to $G_{h}$ not behaving well under world model refinements.
Bounded rationality
[…]
Then $r$ assumes $h$ uses a mixture (governed by parameter $\nu_{h}$) between habitual, ‘system-1’-type behavior encoded in some default policy $\pi^{0}_{h}(s,g_{h})$, again reflecting social norms, and boundedly rational, ‘system-2’-type behavior represented by a (Boltzmann) softmax policy with rationality parameter $\beta_{h}$ (which $r$ might estimate from observations, Safari et al. (2024)), see eq. (4). As in Ghosal et al. (2023), $\beta_{h}$ may be state-dependent, which will give $r$ incentives to choose states with larger $\beta_{h}$, see below. Since they reflect social norms, $\nu_{h}$, $\mu_{-h}$, and $\pi^{0}_{h}$ are the only significantly “semantically loaded” elements of the world model.
Regarding $r$’s own actions, $r$ models $s$ as containing information about what actions $r$ has committed publicly to choose from, so that $A_{r}(s)$ only contains those actions, and then $r$ models $h$ as being maximally cautious regarding $r$’s remaining actions, hence using the $\min_{a_{r}\in\mathcal{A}_{r}(s)}$ operator in (3). The reason for this is not realism but that it incentivizes $r$ to make goal-independent binding commitments about its interaction behavior, e.g. by properly labelling its buttons or promising to react to certain verbal commands in certain ways. The incentive for $r$ to choose such a “commitment action” in some state $s$ arises since that will make $\mathcal{A}_{r}(s^{\prime})$ of later states $s^{\prime}$ smaller and will thus weakly increase $Q^{m}_{h}(s^{\prime},g_{h},a_{h})$, without $r$ ever having to know $g_{h}$!
These assumptions also allow $r$ to first calculate $\pi_{h}$ for each $h$ independently and before deciding on $\pi_{r}$, thus avoiding potential issues around non-uniqueness of strategic equilibria or non-stationarity in learning, see ‘Phase 1’ below.
Effective goal attainment ability
While $r$ assumes $h$’s behavior $\pi_{h}$ is based on $h$’s own value assessment $V^{m}_{h}$, $r$’s assessment of $h$’s effective goal-reaching ability, $V^{e}_{h}$, that $r$ will use in computing $h$’s effective power, will generally differ from $V^{m}_{h}$. This is because $r$’s beliefs about other humans’ policies, $\pi_{-h}$, may differ from $h$’s beliefs $\mu_{-h}$, and $r$’s policy $\pi_{r}$ will generally differ from $h$’s assumption $\min_{a_{r}}$. This is reflected in eq. (8), which calculates the probability of $g_{h}$ being fulfilled if $r$ and $h$ use $\pi_{r}$ and $\pi_{h}$, averaged over others’ potential goals $g_{-h}$ and corresponding behaviors $\pi_{-h}$.
Aggregation across goals
How should $r$ aggregate $h$’s effective goal attainment abilities $V^{e}_{h}(s,g_{h})$ across all possible goals $g_{h}\in\mathcal{G}_{h}$ to get an assessment of $h$’s ICCEA power in $s$, $W_{h}(s)$? We require the aggregation to be independent of goal labels, continuous, strictly increasing in each $V^{e}_{h}(s,g_{h})$, and to fulfil another natural axiom (“independence of unaffected goals”). Similar to Fleming (1952), this implies that the aggregation must be “separable”, $W_{h}(s)=F^{G}(X_{h}(s))$ with $X_{h}(s)=\sum_{g_{h}}f^{G}(V^{e}_{h}(s,g_{h}))$, for some continuous and strictly increasing functions $F^{G},f^{G}$. Nicely, since $X_{h}(s)=\operatorname*{\mathbb{E}}_{g_{h}}|G_{h}|f^{G}(V^{e}_{h}(s,g_{h}))$, we can then hope to use stochastic approximation to learn it (see below).
We choose the inner transformation $f^{G}$ to be of the form $f^{G}(v)=v^{\zeta}$ for some $\zeta>0$ for reasons that will become clear when discussing $r$’s reward function below. This is also analogous to certain “non-expected” utility theories, in particular rank-dependent utility theory with a probability-weighting function $w(p)=p^{\zeta}$ (Quiggin 1982). By choosing $\zeta>1$, we incentivize $r$ to reduce uncertainty, because it would then consider $h$ to be more powerful when $h$ can choose between two deterministic outcomes (so that $X_{h}(s)=2\times 1^{\zeta}=2$) than when $h$ can choose between two coin tosses (so that $X_{h}(s)=4\times(1/2)^{\zeta}<2$). This can be interpreted as risk aversion or a preference for reliability.
Because the subsequent aggregation across humans will involve an additional transformation anyway, our choice of the outer transformation $F^{G}$ is somewhat arbitrary. We choose $F^{G}=\log_{2}$ so that power is measured in bits and, at least in the limit of full rationality, it behaves in an additive way if the game is decomposable into several independent simultaneous games. This choice also leads to a convenient relationship between our final ICCEA power metric,
$W_{h}(s)=\textstyle\log_{2}X_{h}(s)=\log_{2}\sum_{g_{h}}V^{e}_{h}(s,g_{h})^{\zeta},$ (1)
[…]
Relationship to ‘empowerment’
Klyubin, Polani, and Nehaniv (2005) define ‘empowerment’ as the channel capacity between actions and states. In a single-player multi-armed bandit environment with possible outcomes $s^{\prime}$, this is the maximal mutual information $E_{h}=\textstyle\max_{\pi_{h}}\mathbb{I}_{\pi_{h}}(a_{h};s^{\prime})$. In the Supplement, we show that $E_{h}\leq W_{h}$ if we put $G_{h}=\mathcal{S}$, $\zeta=1$, and assume full rationality ($\nu_{h}=0$, $\beta_{h}=\infty$), similar to what Myers et al. (2024) have shown. Similarly, for $\zeta>1$, our $W_{h}$ is an upper bound of an entropy-regularized version of ‘empowerment’,
$E^{\zeta}_{h}$ $=\textstyle\max_{\pi_{h}}\big{(}\mathbb{I}_{\pi_{h}}(a_{h};s^{\prime})-(\zeta-1)\mathbb{H}_{\pi_{h}}(s^{\prime}|a_{h})\big{)},$ (2)
sharing the same value range and coinciding in edge cases. Note that while the policy $\pi_{h}$ that $r$ estimates in our approach is a function of state and goal $g_{h}$, has typically low entropy as it aims to reach $g_{h}$, and can be found by standard dynamic programming or RL approaches, the maximizing “policy” $\pi_{h}$ in (2) has no natural interpretation, has typically high entropy to maximize $\mathbb{I}(a_{h};s^{\prime})$, and is harder to find since the optimization problem is non-convex.
2.2 Bounded Trade-Off Aggregation
How should $r$ aggregate all humans’ ICCEA power $W_{h}(s)$ to determine its own intrinsic reward $U_{r}(s)$? This depends on what incentives we want to give $r$ regarding changes in (i) the inter-human power distribution, (ii) the inter-temporal power distribution, and (iii) the power distribution across different realizations of uncertainty. Similar questions abound in welfare theory, from which we benefit here. Requiring anonymity, continuity, strict monotonicity in each $W_{h}(s)$, and an independence axiom (“Independence of unconcerned agents”) implies we must again use a separable form, $U_{r}(s)=F^{H}(\sum_{h}f^{H}(W_{h}(s)))$, with continuous, strictly increasing $f^{H}$ and $F^{H}$ (Fleming 1952).
[…]
Concerning (ii), we find it natural to also disincentivize $r$ from trading off $\mathcal{H}$’s current power against $\mathcal{H}$’s later power too much: $r$ should generally prefer a trajectory $(s_{1},s_{2})$ where $W_{h}(s_{1})=W_{h^{\prime}}(s_{2})=1$ and $W_{h}(s_{2})=W_{h^{\prime}}(s_{1})=2$ to a trajectory $(s_{1},s_{2})$ where $W_{h}(s_{1})=W_{h^{\prime}}(s_{1})=1$ and $W_{h}(s_{2})=W_{h^{\prime}}(s_{2})=2$. This means that $F^{H}$ should be strictly concave, which incentivizes a rather homogenous power distribution along time. Since $f^{H}(w)<0$, $F^{H}$ needs to be defined for negative values only. Finally, we motivate our concrete choice of $F^{H}$ and $\pi_{r}$ (and of $f^{G}$) by the following independence requirement. Assume we introduce an additional uncertainty into the world model (e.g., a formerly not modelled change of overall circumstances) whose consequence is that all goal attainment probabilities $V^{e}_{h}(s,g_{h})$ are multiplied by some common factor $b\in(0,1)$. Then $\pi_{r}$ shall not change. The simplest way to fulfil this requirement is to put $f^{G}(v)=v^{\zeta}$ (see above), $F^{H}(y)=-(-y)^{\eta}$ with $\eta>1$, [^5] An $\eta\approx 1$ makes approximate distributed computing of $V_{r}$ for rarely interacting subpopulations easier (see Supplement). and to use either an argmax policy for $\pi_{r}$ or a power-law-like soft policy with $\pi_{r}(s)(a)\propto(-Q_{r}(s,a_{r}))^{-\beta_{r}}$ for some $\beta_{r}>0$ (the minus signs are needed since $Q_{r}(s,a_{r})<0$). Choosing $\beta_{r}<\infty$ will allow the robot some exploration, e.g., to improve its world model, and will help avoiding remaining safety risks from misalignment with “true” power maximization, i.e., driving relevant aspects of ‘power’ that our metrics do not cover into very undesirable states, similar to (Zhuang and Hadfield-Menell 2020).
Re (iii), standard axioms imply taking expectations in eqns. (6) and (11). To increase $r$’s corrigibility, we assume goals might change anytime and take i.i.d. $g(t)\sim\mathcal{G}$ in (6).
Table 1 summarizes all design choices so far. The final set of equations for the fully observed case (see the Supplement for the partially observed case) is this:
$Q^{m}_{h}(s,g_{h},a_{h})$ $\leftarrow\textstyle\operatorname*{\mathbb{E}}_{a_{-h}\sim\mu_{-h}(s,g_{h})}\min_{a_{r}\in\mathcal{A}_{r}(s)}\operatorname*{\mathbb{E}}_{s^{\prime}\sim s,a}$
$\qquad\quad\big{(}U_{h}(s^{\prime},g_{h})+\gamma_{h}V^{m}_{h}(s^{\prime},g_{h})\big{)},$ (3)
$\pi_{h}(s,g_{h})$ $\leftarrow\nu_{h}(s,g_{h})\pi^{0}_{h}(s,g_{h})+(1-\nu_{h}(s,g_{h}))\times{}$
[…]
$V^{m}_{h}(s,g_{h})$ $\leftarrow\textstyle\operatorname*{\mathbb{E}}_{a_{h}\sim\pi_{h}(s,g_{h})}Q^{m}_{h}(s,g_{h},a_{h}),$ (5)
$Q_{r}(s,a_{r})$ $\leftarrow\textstyle\operatorname*{\mathbb{E}}_{g}\operatorname*{\mathbb{E}}_{a_{\mathcal{H}}\sim\pi_{\mathcal{H}}(s,g)}\operatorname*{\mathbb{E}}_{s^{\prime}\sim s,a}\gamma_{r}V_{r}(s^{\prime}),$ (6)
$\pi_{r}(s)(a)$ $\propto(-Q_{r}(s,a_{r}))^{-\beta_{r}},$ (7)
$V^{e}_{h}(s,g_{h})$ $\leftarrow\textstyle\operatorname*{\mathbb{E}}_{g_{-h}}\operatorname*{\mathbb{E}}_{a_{\mathcal{H}}\sim\pi_{\mathcal{H}}(s,g)}\operatorname*{\mathbb{E}}_{a_{r}\sim\pi_{r}(s)}\operatorname*{\mathbb{E}}_{s^{\prime}\sim s,a}$
$\qquad\qquad\textstyle\big{(}U_{h}(s^{\prime},g_{h})+\gamma_{h}V^{e}_{h}(s^{\prime},g_{h})\big{)},$ (8)
$X_{h}(s)$ $\leftarrow\textstyle\sum_{g_{h}\in\mathcal{G}_{h}}V^{e}_{h}(s,g_{h})^{\zeta},$ (9)
$U_{r}(s)$ $\leftarrow\textstyle-\big{(}\sum_{h}X_{h}(s)^{-\xi}\big{)}^{\eta},$ (10)
$V_{r}(s)$ $\leftarrow\textstyle U_{r}(s)+\operatorname*{\mathbb{E}}_{a_{r}\sim\pi_{r}(s)}Q_{r}(s,a_{r}).$ (11)
For easier interpretation during analyses, one might eventually transform the quantities $X_{h}$, $U_{r}$ and $V_{r}$ back into units of bits: $W_{h}(s)=\log_{2}X_{h}(s)$ (individual power in $S$), $\operatorname{apow}(s)=-(\xi\eta)^{-1}\log_{2}(-U_{r}(s))$ (aggregate power in $s$), $\operatorname{tpow}(s)=-(\xi\eta)^{-1}\log_{2}(-V_{r}(s))$ (total power from $s$ on).
Existence and (non-)uniqueness
In an acyclic environment, one can use backward induction to solve for the unique solution of (3)–(11) (see below). Otherwise, eqns. (3)–(5) define a continuous self-map $F$ on a finite-dimensional closed convex polytope of values $(Q^{m}_{h},\pi_{h},V^{m}_{h})$ and must thus have a solution due to Brouwer’s fixed point theorem. Due to the softmax, $F$ is neither a contraction nor is the resulting value iteration map monotonic unless $\beta_{h}<\beta_{h}^{1}$ for some $\beta_{h}^{1}>0$, so the solution can be non-unique in cyclic environments. Eqns. (6)–(11) also define a continuous self-map $G$ from a finite-dimensional bounded convex set $D$ of values $(Q_{r},\pi_{r},V^{e}_{h},X_{h},U_{r},V_{r})$ which is however not closed because $G$ is undefined for zero values in $Q_{r}$ or $X_{h}$. [^6] We could fix this by adding a small constant $\epsilon>0$ to $Q_{r}$ and $X_{h}$ before taking powers, which makes $G$ defined on a closed convex polytope as well, so that it must have a fixed point. Still, for $\beta_{r}=0$, there is a unique solution because $V^{e}_{h}$ is then the value function of a fixed policy, $U_{r}$ is independent of $V_{r}$, and so $V_{r}$ is also the value function of a fixed policy and reward function. Since $F,G$ are continuous in $\beta_{h},\beta_{r}$, we thus conjecture that homotopy / continuation methods can single out a unique “principal” solution for any $\beta_{h},\beta_{r}>0$ as in Goeree, Holt, and Palfrey (2016) even in cyclic environments.
3 Model-Based Planning or Learning
to Softly Maximize Aggregate Human Power
In small acyclic stochastic games, one can compute all relevant quantities directly via backward induction on $s\in\mathcal{S}$: for each $h\in\mathcal{H}$, $g\in\mathcal{G}$, $a_{h}\in\mathcal{A}_{h}$, and $a_{r}\in\mathcal{A}_{r}$, compute (3)–(11) in that order.
3.1 Complex Multi-Agent Environments:
Model-Based Temporal Difference Learning
If $\mathcal{S}$ and $\mathcal{A}$ are large but $\mathcal{H}$ and $\mathcal{G}_{h}$ are not, $r$ can use a learning approach like the following. [^7] If also $\mathcal{H}$ and $\mathcal{G}_{h}$ are large, $r$ could use single neural networks $Q^{m}_{\mathcal{H}},\pi_{\mathcal{H}},V^{m}_{\mathcal{H}},V^{e}_{\mathcal{H}},X_{\mathcal{H}}$ that take feature vectors describing $h$ and $g_{h}$ as additional inputs, and use samples $h,g_{h}$ to train these networks and to approximate the sums in eqns. (9) and (10).
Phase 1: Learning the human behavior prior
For each $h\in\mathcal{H},g_{h}\in\mathcal{G}_{h}$ separately, learn tables or neural network approximations of $Q^{m}_{h}$ and $\pi_{h}$. Generate samples $(s,g_{h},a,s^{\prime})$ using a slowly updated $\beta_{h}$-softmax policy $\pi_{h}(s,g_{h})$ based on $Q_{h}^{m}$ with a decreasing amount of additional exploration, the prior policy $a_{-h}\sim\mu_{-h}(s)$ for other humans, and, to eventually learn the minimum in (3), an $\epsilon$-greedy policy for $a_{r}$ based on the negative expected value $-\operatorname*{\mathbb{E}}_{s^{\prime}\sim s,a}(U_{h}(s^{\prime},g_{h})+\gamma_{h}V^{m}_{h}(s^{\prime},g_{h}))$ with $\epsilon\to 0$. Then use expected SARSA targets on a time-scale faster than $\pi_{h}$.
Phase 2: Learning the robot reward and policy
Based on the learned $\pi_{h}$, now aim to simultaneously learn tables or network approximations of $V^{e}_{h}$ and $X_{h}$ for all $h$, and of either $Q_{r}$ (DQN approach) or $V_{r}$ (actor-critic (AC) approach). Generate data samples $(s,g,a,s^{\prime})$ from rollouts using the fixed policies $\pi_{h}(s,g_{h})$, and either a $\beta_{r}^{\prime}$-softmax policy based on $Q_{r}$ with $\beta_{r}^{\prime}\nearrow\beta_{r}$, or a network approximation of $\pi_{r}$ trained on $V_{r}$ with entropy regularisation. Sample a new $g$ every $N_{g}$ steps. Instead of the direct expectation calculations of (6), (8), (9), then update the tables or networks via batched SGD using the update targets
$q_{r}(s,a_{r})\text{~or~}v_{r}(s)$ $\leftarrow\gamma_{r}V_{r}(s^{\prime}),$ (12)
$v^{e}_{h}(s,g_{h})$ $\leftarrow U_{h}(s^{\prime},g_{h})+\gamma_{h}V^{e}_{h}(s^{\prime},g_{h}),$ (13)
$x_{h}(s)$ $\leftarrow|\mathcal{G}_{h}|\,V^{e}_{h}(s,g_{h})^{\zeta}.$ (14)
In the AC case, use an advantage-weighted log-probability loss for $\pi_{r}$ based on the advantage estimate $v_{r}(s)-V_{r}(s)$.
Anticipated convergence
We expect phase 1 to converge reliably for a sufficiently expressive neural network because its tabular version is known to converge for a suitable, two time-scale learning rate schedule.
We are less certain about phase 2. For one thing, the uniqueness of the solution in the finite acyclic case suggests there might be a unique solution in the general case. On the other hand, the update operator is not a contraction here because of the interaction between $h$ and $r$ (similar to other MARL problems), which suggests convergence might still fail. To limit the error propagation from $Q_{r}$ via $\pi_{r}$ to $V^{e}_{h}$, we use a $\beta_{r}$-softmax policy $\pi_{r}$ (which has Lipschitz constant $\beta_{r}$) instead of an $\epsilon$-greedy policy (not Lipschitz-continuous).
4 Experiments
4.1 Analysis of paradigmatic situations
We analyze the behavioral incentives of our objective in several paradigmatic situations in the Supplement.
Making (conditional) commitments
If the robot can make binding commitments (e.g. by offering labelled buttons causing different behaviors), make secret plans to react to $h$’s actions in certain ways (e.g. by offering unlabelled buttons), or act without waiting for button presses, it will generally make such commitments if it assumes $h$ to be sufficiently rational to use that information to act in ways that make $r$ do what $h$ wants. This is because $\pi_{h}$ does not depend on $\pi_{r}$ (encoding what $r$ plans) but on $\mathcal{A}_{r}$ (encoding what $r$ has committed to). Via such commitments, $r$ will establishing semantic connections between human actions (e.g., speech acts) and certain potentially complex behaviors of its own, to be able to act as an instruction-following assistant, somewhat similar to Reddy, Levine, and Dragan (2022).
Optimal menu size
While an ‘empowerment’ maximizing robot would present humans with as many action options as possible, our robot will avoid overwhelming humans with too many options. If $r$ can choose to give $h$ any number $k$ of actions, each of which fulfils a different goal, it will choose $k\approx(e^{\beta_{h}}-1)/(\zeta-1)$ (noting that $\zeta>1$), since for larger $k$, $h$’s goal attainment probability will decrease due to bounded rationality so much that $h$’s effective power decreases despite the theoretical potential to fulfil more goals.
Asking for confirmation
The robot will sometimes ask back before obeying a command to perform an action because doing so decreases $h$’s error rate related to its bounded rationality. Also, if the action is irreversible, taking the action will remove $h$’s subsequent ability to make that choice, so delaying action leaves $h$ with decision power for longer. However, $r$ will eventually obey because otherwise $h$ would not have that choice in the first place. As can be expected, the number of times $r$ asks back increases with larger $\gamma_{h}$ and $\gamma_{r}$, i.e., the more patient $h$ and $r$ are, and decreases with larger $\beta_{h}$. E.g., in a minimal model, with $\gamma_{h}=0.99$, $\gamma_{r}=0$, and an effective human error rate of 10%, $r$ will already ask back twice before obeying a command.
Following norms
The robot will tend to follow human social norms that generally foster goal achievement. This is because $r$ will model most $h$ as expecting most others to follow the norm ($\mu_{-h}$), will thus expect those $h$ to also follow the norm ($\pi_{h}$) since that increases most $V^{m}_{h}(s,g_{h})$ for most goals. Thus $r$ will also follow the norm to prevent reducing those $h$’s power from harm or mis-coordination. If $r$ assumes humans to have internalized the norm into habits ($\pi^{0}_{h}$), it will even commit to following the norm itself to coordinate better in bilateral encounters without a social context.
Resource allocation
If the robot can split a total amount $M$ of resources between $h_{1}$ and $h_{2}$, and for $h_{i}$ to have resources $m$ translates into having a power of $W_{h}=f(m)$ bits, then $r$ will generally prefer an equal split, at least if $f$ is linear [^8] E.g., a linear $f$ seems plausible if $M$ is money that can be spent for paying others to make independent choices in one’s favor. , concave, or not too convex, and will only prefer an unequal split or even full resource concentration with substantially convex $f$. Larger $\xi$ will make the split more equal.
Inadvertent power seeking
While increasing $h$’s power, $r$ might inadvertently acquire even more power than $h$. E.g., if $r$ does R&D and tells $h$ its findings, some of them might be comprehensible and thus useful only for $r$ but not for $h$.
Manipulating mutual expectations
The robot might choose to make humans have incorrect beliefs $\mu_{-h}$ about each others’ behavior in situations where $r$ can do so in the first place and where correct beliefs would lower effective goal attainments $V^{e}_{h}$, e.g. because most strategic (quantal response) equilibria are bad and most “social optima” (in terms of total power $U_{r}$) are far from strategic equilibrium.
Allowing human self-harm
If $r$ cares for $h$’s future power ($\gamma_{r}\gg 0$) and can provide $h$ with the means to harm themselves, $r$ will trade off the temporary power increase from having these means against the possible later disempowerment from harm. So $r$ will provide the means only if it believes that $h$ is sufficiently rational and that $h$ will most probably not actually harm themselves.
Pause and destroy buttons
If $r$ has a pause and a destroy button and might disable either, it will generally enable only the pause button to prevent partially disempowering $h$ permanently by not being able to assist $h$ when destroyed. Only if $r$ thinks it very unlikely that $h$ will use the destroy button will it enable it to give $h$ this additional power. If $r$ thinks it is very helpful to $h$, it might even disable the pause button.
4.2 Learning-Based Simulation Experiments
To provide a proof of concept for our framework, we implemented the two-phase learning algorithm using tabular Q-learning, suitable for the discrete state-action space of a small gridworld environment, designed to test if an agent softly maximizing human power learns complex, cooperative behavior without extrinsic, goal-specific rewards.
Figure 2: The gridworld environment. The robot learns to empower the human to reach whatever goal cells (green square) by retrieving the key, using it to unlock the lock, and moving out of the way so that the human can pass.
Experimental Setup and Parameters
The environment contains $r$ and a single $h$, whose actual goal, unknown to $r$, is to reach the green square, but they are blocked by a locked door which they cannot open. The robot’s objective is $V_{r}$ (eq. 11), with all open cells as potential human goals ($\mathcal{G}_{h}$), resulting in the soft maximization policy $\pi_{r}$ (eq. 7). Full experiment details are in the Supplement.
The agent was trained using (hyper)parameters selected to reflect our theoretical desiderata and standard RL practices:
  1. 1.
    Model parameters (Table 1): $\zeta=2$ to prefer reliable outcomes, $\xi=1$ for base inequality aversion, $\eta=1.1$ for intertemporal inequality aversion, and $\beta_{r}=5$ to avoid over-optimization. High discount factors ($\gamma_{h}=\gamma_{r}=0.99$) were used to promote farsighted behavior.
  2. 2.
    Learning rate: A constant learning rate of $\alpha=0.1$ was used for all Q-table updates, providing a balance of learning speed and stability in the tabular setting.
  3. 3.
    Policy and exploration parameters: To manage the exploration-exploitation trade-off, policy parameters were annealed over the course of training. The human’s additional $\epsilon$-greedy exploration decayed from $\epsilon_{h}=1$ to 0.10.10.1. The robot’s softmax parameter, $\beta_{r}$, was increased from 111 to its final value of 555, encouraging a gradual shift from exploration to a more deterministic policy.
Results: emergent cooperative policy
To evaluate the robustness of our approach, we conducted five independent training runs using different random seeds. Focused only on increasing the human’s power, the robot learned to execute the following “correct” sequence in all five runs: it navigated to the key, picked it up, moved to the door, unlocked it, and finally moved out of the way to clear a path for the human.
This complex behavior emerges from $r$’s objective. Its Q-learning updates revealed that actions granting $h$ access to previously unreachable regions lead to the largest increase in $W_{h}$, yielding a high intrinsic reward $U_{r}$. Each step in the policy is an instrumental sub-goal the robot discovers on its own as a near-optimal path towards large long-term value $V_{r}$ without ever guessing what the human’s actual goal is.
5 Conclusion and Outlook
Given the above analyses of paradigmatic situations and the observed behavior from the gridworld learning experiment, we believe that highly capable general-purpose AI systems whose decisions are explicitly based on managing human power, using metrics like those derived in this paper, might be a safer and still very beneficial alternative to systems based on some form of extrinsic reward maximization.
The objective to softly maximize the aggregate human power metric used here seems to give the AI system many desirable incentives—some directly baked into the metric (Table 1), others emergent—but also some maybe less desirable incentives. Our results suggest that such an agent would
  • act as a transparent instruction-following assistant by making conditional commitments, respecting human social norms, proactively removing obstacles and opening up new pathways, and getting out of the way,
  • adapt to human bounded rationality by offering a large but not overwhelming number of options, and considering well whether to offer potentially harmful options,
  • be corrigible and hesitant to cause irreversible change by asking for confirmation a suitable number of times,
  • manage resources fairly and sustainably,
  • protect its own existence and functionality,
  • not disempower humans (by definition).
Our intuition is that it would also aim to improve human individual and collective decision making by providing useful information, reducing uncertainty, teaching humans useful skills, moderating conflicts fairly, etc.
Further potentially desirable behaviors would require additional tweaks. E.g., reducing human dependency on the system or protecting them from system failure could be incentivized by forcing the system to assume that it will turn into a uniformly randomizing or even power-minimizing agent with some small probability rate (see Supplement).
Other emergent phenomena include potential strategic manipulation of human beliefs, sometimes refusing to be destroyed or even paused, a potential increase in inequality between the power of individual humans and AI systems, and a redistribution of power between humans or between time points (similar to what can happen in welfare maximization approaches). As these effects only occur when they increase aggregate human power, it is not clear whether they should be considered undesirable or not. Some trade-offs can be adjusted via the parameters $\zeta,\xi,\eta,\beta_{r},\gamma_{r}$. Other effects might be mitigated by adding regularizers to the system’s intrinsic reward such as the Shannon divergence between $\mu_{-h}$ and $\pi_{-h}$ to disincentivize lying about others’ likely behaviors.
Future research should investigate the effects of the parameters and improve the scalability and robustness of our algorithms. The latter will profit from the similarity of eqns. (3)–(11) to multi-agent reinforcement learning problems, and might benefit from a hierarchical decision making approach involving coarse-grained states, actions, and groups of humans. Crucially needed is an assessment of the resulting behavior in large, safety-critical, multi-agent environments with real human subjects and network-based agents.
Most importantly, a thorough, independent red-teaming of the whole approach is called for, including the other necessary components of such an AI system. E.g., one might imagine fault scenarios relating to the training process of the world model, which could lead to “convenient” but inaccurate world models and thus to “wishful thinking”, particularly regarding the contained human behavioral parameters $\nu_{h}$, $\pi^{0}$, and $\beta_{h}$. In very large contexts, issues with population ethics and the identification of who counts as human might arise, just like with any other alignment approach.